jaypee calculated his total revenue for the night to be P54,540. a total of 540 tickets and souveneirs were sold. If a ticket sold at P112 each and the souvenier cost P68, how much tickets were sold?
Answer by Boreal(15235)

T+S=540
112T+68S=54540
multiply top by -68 to eliminate S
-68T-68S=-36720
44T=17,820
T=405
pick 405 tickets@112=45360
then 135 souvenirs@68=9180, total is 54540